OpenStudy (anonymous):

The x-intercept of a line is -5 and the y-intercept of the line is -2. What is the equation of the line? Y=-5/2x-5 Y=2/5x+2 Y=5/2x-2 Y=-2/5x-2

OpenStudy (anonymous):

Also, please explain how you got the answer

OpenStudy (azureilai):

d. you can graph it, then count the slope between the x and y intercepts and then plug in the y intercept at the end

OpenStudy (anonymous):

the y-intercept appears in the equation since they are in slope-intercept form where y=mx+b b=y-intercept so we answers are narrowed to C and D no but y=0 to each equation to see which gives you x=-5

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Y=5/2x-2 Y=-2/5x-2 0=5/2x-2 0=2/5x-2 2=5/2x 2=2/5x x=4/5 x=5 so the answer would be D